# figureb tableb grand_reference.bib

### v1

@@ I Introduction
The K-Means, one of the top 10 algorithms in data mining [wu2008top], is widely used in image classification, vector quantization [gersho2012vector], knowledge discovery [fayyad1996advances], and pattern classification [duda1973pattern]. However, K-Means is increasingly vulnerable to transient faults caused by high circuit density, low near-threshold voltage, and low near-threshold voltage [lutz1993analyzing, nicolaidis1999time, laprie1985dependable]. Oliveira et al. [oliveira2017experimental] demonstrated an exascale system with 190,000 cutting-edge Xeon Phi processors that still suffer from daily transient errors under ECC protection. Recognizing the importance of this issue, the U.S. Department of Energy has named reliability as a major challenge for exascale computing [lucas2014doe].
@@
Intel Corporation first documented a transient error leading to soft data corruption in 1978, marking a significant recognition of the impact of such faults in both academia and industry [may1979alpha]. Subsequently, in 2000, Sun Microsystems reported server crashes and outages at major sites like America Online and eBay back to cosmic ray strikes on unprotected caches [baumann2002soft]. Similarly, Virginia Tech, in 2003, had to dismantle and sell its newly assembled Big Mac cluster of 1100 Apple Power Mac G5 computers due to a lack of ECC protection that resulting the system prone to cosmic ray-induced failures [geist2016supercomputing]. Despite advancements in ECC protection, transient faults continue to challenge system reliability. For example, a simulation by Oliveira et al. of an exascale system with 190,000 advanced Xeon Phi processors revealed daily vulnerabilities to transient errors, even with ECC [oliveira2017experimental]. These faults are not just theoretical concerns; real-world impacts have been recorded by Google, which experienced transient faults causing incorrect outputs in its production environment [hochschild2021cores]. In 2018, faced with the ongoing risk of transient faults on its large-scale infrastructure, Meta launched an internal investigation to seek solutions [dixit2021silent]. Transient faults may cause either fail-stop errors, which lead to system crashes, or fail-continue errors, which produce incorrect outcomes. While fail-stop errors can often be addressed through checkpoint/restart mechanisms [phillips2005scalable, NEURIPS2019_9015, tao2018improving, tensorflow2015-whitepaper] or algorithmic approaches [chen2009optimal, wu2011fault, hakkarinen2012multilevel, hakkarinen2014fail, chen2008scalable, chen2008extending], fail-continue errors are more problematic as they silently corrupt the state of applications and result in incorrect outputs [mitra2014resilience, cher2014understanding, dongarra2011international, calhoun2017towards, snir2014addressing]. These types of errors are particularly dangerous in environments where safety is critical [li2017understanding]. This paper focuses on fail-continue errors within the computational logic units, assuming that errors in memory and those causing system stops are managed using error-correcting codes and checkpoint/restart techniques. We refer to these issues as soft errors.
@@
Existing fault tolerance methods are feasible for K-Means clustering. Taamneh [taamneh2020parallel] proposes a checkpointing strategy, which involves periodically saving the centroids to stable storage during normal operation and restarting from checkpointed centroids in the event of a failure. However, this method cannot detect silent errors and requires recomputation after an error. Dual modular redundancy (DMR) verifies computational correctness by replicating instructions and comparing the results [oh2002error, oh2002control, reis2005swift, yu2009esoftcheck, chen2016simd]. Figure 1 left illustrates the K-Means workflow. We find that DMR can protect the memory-bound routine efficiently, such as the centroids update phase in Figure 1 left step 3. It is because the memory latency of loading the data points is so high that all arithmetic operations can be duplicated without introducing an overhead over $1\%$, even for synchronous instructions like atomicAdd. However, DMR failed to protect the compute-bound distance calculation in Figure 1 left step 1 due to the redundant computations. In Figure 1 left step 2, nearest cluster matching is fused within the distance computation kernel to eliminate redundant memory acess. Algorithm-based fault tolerance (ABFT) reduces this redundancy by using checksums based on equivalence relationships, lowering the redundancy to $O(1/N)$, where $N$ is the problem size [chen2008algorithm]. Efforts have been made to apply ABFT on GPUs, designing kernel fusion schemes to minimize the overhead of ABFT by hiding it within the memory transactions and computing unit gaps [ding2011matrix, zhai2021ft, kosaian2021arithmetic, wu2023anatomy].
@@
However, existing ABFT fusion strategies lack architecture-aware optimization. Kosaian [kosaian2021arithmetic] proposes an error detection scheme for tensor-core but the error correction requires a time-redundant recomputation. Although error correction is discussed in [zhai2021ft, wu2023anatomy], they only utilize a register-reusing which becomes outdated for Amepere architecture. The register reusing stage is illustrated in the SIMT GEMM part of Figure 1 and the asynchronous copy is presented in the Tensor-core part. Before Ampere architecture, the global-to-share memory transfer passes through the register file explicitly, enabling register reusing for checksum computation. However, the latest asynchronous memory copy enables a new data path from global to shared memory bypassing the register. Without register reusing, checksum computation requires additional expensive memory read, destroying the carefully designed pipeline. Additionally, these fault tolerance strategies consider matrix multiplication individually, with no aware of the underlying K-Means routine.
@@
Furthermore, the state-of-the-art K-Means implementation, cuML suffers from fixed kernel parameters, resulting in sub-optimal and low utilization of peak computing performance over various problem sizes. For example, the K-Means in cuML obtain a performance of less than 10% of the peak performance for both single and double precisions. Due to the underlying tall-and-skinny matrix multiplication computation inside the K-Means, a code generation strategy and a parameter selection scheme are necessary to improve the K-Means performance for various of input shapes and data types.
@@
To address these issues, we propose FT K-Means, a high-performance K-Means implementation equipped with an algorithm-based fault tolerance scheme that detects and corrects silent data corruptions at computing units on-the-fly. More specifically, our contributions include the following:
@@
  • Experimental results of single precision and double precision on an NVIDIA A100 server GPU and a Tesla Turing T4 GPU show that FT K-Means offers a $10\%-300\%$ improvement compared to the state-of-the-art library cuML. The fault tolerance scheme in FT K-Means introduces an average overhead of 11%, even under tens of error injections per minut.
@@ II-A Fault Model
FT K-Means is designed to identify and correct errors in computing units that could influence the final results. It operates under the assumption that memory errors are managed by ECC [bird2017neutron] and communication reliability issues are addressed by FT-MPI [fagg2000ft]. For compute errors during runtime, a fault-tolerant strategy is implemented based on the single-event upset (SEU) assumption [binder1975satellite, petersen2013single, binder1975satellite], which posits that only one soft error occurs within each detection and correction interval. This assumption is supported by the low frequency of multiple soft errors, attributed to brief fault detection intervals—a concept prevalent in several studies [reis2005swift, zhai2021ft, ding2011matrix, wu2014ft]. The fault tolerance method involves high reliability in detecting faults with minimal false alarms [turmon2000software]. Specifically, each threadblock randomly selects an element to corrupt by flipping a single bit, either in its 32-bit float representation or 64-bit double representation. A checksum test with a defined threshold $\delta$ then attempts to operate the corrupted computations, followed by an application of an error correction scheme.
@@ II-C Algorithm-Based Fault Tolerance
Soft error protection algorithms aim to identify and rectify errors that can arise in iterative or computationally demanding applications, such as signal processing [johnston2023curriculum, liu2023stationary, johnston2023downlink] and scientific computing [liu2023cusz, huang2023exploring, jian2024cliz, liu2024high]. The development of these algorithms dates back to 1984, with Huang [huang1984algorithm] first specifically designed for matrix-matrix multiplication. The fundamental concept of these algorithms involves encoding matrices $X$ and $Y$ into checksums $X^{c}$ and $Y^{r}$, respectively. This encoding is achieved through the following equations.
@@
where, $i$ represents the step number in the outer-product update of the matrix $D$, with $D_{i}$ indicating the result of each outer-product, namely $X^{\prime}(:,i)\cdot Y^{\prime}(i,:)$. The offline version of the double-checksum approach can correct only one error in the entire computation. In contrast, the online version corrects a single error at each step of the update, allowing it to address multiple errors throughout the entire program. Ding et al. [ding2011matrix] first presented an implementation of the outer-product ABFT-GEMM for GPUs. To minimize the memory latency from checksum operations, Zhai et al. developed combined compute kernels for GEMM on AVX-512-enabled CPUs [zhai2021ft, zhai2023ft]. Kosaian and Rashmi [kosaian2021arithmetic] present a warp-level ABFT implementation capable of error detection, but not correction. Shixun et al. [wu2023anatomy] proposed a fully-fused ABFT-GEMM that effectively detects and corrects computational errors. However, the kernel fusion strategy relies on reusing registers during transfers between global and shared memory. This approach does not extend well to GPU architectures post-Turing, e.g. Ampere, due to the introduction of asynchronous copy instructions. Besides matrix multiplication, ABFT scheme is widely adopted in fast Fourier transforms [liang2017correcting, wu2024turbofft], sorting [li2019ft], iterative methods [chen2013online, tao2016new, chen2014extending], and computer vision [zhao2020algorithm].
@@ III-A3 Kernel Fusion in thread and thread block level
After GEMM, we write the result matrix back to GPU memory and load the matrix again in order to do a row-wise reduction. This greatly increases the amount of data movement and increases storage overhead. In this part, we apply kernel fusion [zhai2023bytetransformer, zhai2023architectural] to accomplish part of the reduction within the GEMM kernel. Firstly, as Figure 2 Step 2 indicates, each thread in the GEMM kernel handles a small submatrix of the result matrix. After the computation of GEMM, we can simply find the minimum column in each row within this submatrix, and write it to shared memory. After all threads in the threadblock finish this step, thread 0 reads all results in shared memory, calculates another row-wise minimum for each row as the final partial answer for this threadblock, and writes it to GPU memory. Assuming each threadblock has size $TM\times TN$, the reduction kernel only needs to load and handle $TN/N$ of data compared to the last part, which obtains a speedup of 1.13x compared to last part.
@@ III-B Automatic Code Generation
Default naming of kernel parameters A group of kernel parameters in cuML and CUTLASS refers to a set of parameters, threadblock level parameters, warp level parameters and thread level parameters. Each level is composed of three parameters from each dimension. For example, in warp level, the three parameters are labeled Warp.M, Warp.N, and Warp.K, and they refer to the three dimensions of M, N, and K in GEMM.
@@
Code Generation Strategy Figure 3 demonstrates the method we used in our code generation strategy. The code structure of the CUTLASS GEMM kernel in cuML are as the left side of the figure. Firstly, the epilogue is integrated into the FusedDistanceNN kernel. The FusedDistanceNN kernel handles both GEMM and reduction. And then the kernel is wrapped into GemmKernel. With threadblock, warp, and thread level parameters set, it is then transferred to the next template as FusedDistanceNNGemm. Finally, with all K-Means routine sets, the whole cutlassFusedDistanceNN template function works as an interface of the cluster assignment stage. In order to generate a set of feasible kernels with customized parameters, we write a code to test all possible parameters in the search space defined by ourselves, as shown in the codegen part. And for every group of parameters, try it in a demo code. If it can compile and run, which means it is functionally correct. And then we put it in the parameters queue. With all parameters, we then run a code generator to modify the three parts of the original cuML source code, injecting corresponding functions for each parameter in each stage.
@@ III-B1 Kernel Parameters
The kernel parameters used in code generation is not chosen by brute forcing every possible integer in a range for every parameter in the parameter group (a parameter group means: threadblock level, warp level and thread level parameters). We follow some rules. 1) all parameters must be power of 2 2). Warp.K = Threadblock.K. 3). warp size/thread size is 8 or 16. 4). thread size is fixed for FP32 (16, 8, 4) and FP64 (8, 8, 4) owing to the size of the tensor core.
@@ III-B2 Code Generation Template
The code generation strategy defines the K-Means kernel using different tiling parameters. For single precision, 157 kernels with different parameter sets are defined while 145 kernels are defined for double precision. The test workflow illustrated in Figure 3 checks the feasibility of those kernels and performs the benchmark over 64 problem sizes. The benchmark result of different kernels will be employed as the kernel selection criterion. Below, we give a brief explanation of the K-Means kernel so that we can switch to the discussion of fault tolerance part.
@@ V Performance Evaluation
We evaluate our K-means on two NVIDIA GPUs, a Tesla Turing T4 and a 40GB A100-PCIE GPU. The Tesla T4 GPU is connected to a node with two 16-core Intel Xeon Silver 4216 CPUs, whose boost frequency is up to 3.2 GHz. The associated CPU main memory system has a capacity of 512 GB at 2400 MHz. The A100 GPU is connected to a node with one 64-core AMD EPYC 7763 CPU with a boost frequency of 3.5 GHz. We compile programs using CUDA $\mathtt{11.6}$ with the $\mathtt{-O3}$ optimization flag on the Tesla T4 machine, and using CUDA $\mathtt{12.0}$ on the A100 machine. A100 has a peak computational performance of $19.5$ TFLOPS for single precision and $9.7$ TFLOPS for double precision. The memory bandwidth is $1.55$ TB/s. T4 has a peak performance of $8.1$ TFLOPS for single precision and a peak performance of $0.253$ TFLOPS for double precision. The bandwidth of T4 is $320$ GB/s. We first demonstrate the benchmark result between FT K-Means without fault tolerance and cuML for FP32 and FP64 on A100. Next, we evaluate the FT K-Means under fault tolerance. Then, we benchmark FT K-Means under error injections with cuML, Wu’s ABFT, and Kosaian’s ABFT. Finally, we present the same performance evaluation on the T4 GPU. All experimental results are averaged over ten trials.

### v2

@@ I Introduction
The K-means, one of the top 10 algorithms in data mining [wu2008top], is widely used in image classification, vector quantization [gersho2012vector], knowledge discovery [fayyad1996advances], and pattern classification [duda1973pattern]. However, K-means is increasingly vulnerable to transient faults caused by high circuit density, low near-threshold voltage, and low near-threshold voltage [lutz1993analyzing, nicolaidis1999time, laprie1985dependable]. Oliveira et al. [oliveira2017experimental] demonstrated an exascale system with 190,000 cutting-edge Xeon Phi processors still suffering from daily transient errors under ECC protection. Recognizing the importance of this issue, the U.S. Department of Energy has named reliability as a major challenge for exascale computing [lucas2014doe].
@@
Intel Corporation first documented a transient error leading to soft data corruption in 1978, marking a significant recognition of such faults’ impact in academia and industry [may1979alpha]. Subsequently, in 2000, Sun Microsystems reported server crashes and outages at major sites like America Online and eBay back to cosmic ray strikes on unprotected caches [baumann2002soft]. Similarly, Virginia Tech, in 2003, had to dismantle and sell its newly assembled Big Mac cluster of 1100 Apple Power Mac G5 computers due to a lack of ECC protection that resulting the system prone to cosmic ray-induced failures [geist2016supercomputing]. Despite advancements in ECC protection, transient faults continue to challenge system reliability. For example, a simulation by Oliveira et al. of an exascale system with 190,000 advanced Xeon Phi processors revealed daily vulnerabilities to transient errors, even with ECC [oliveira2017experimental]. These faults are not just theoretical concerns; real-world impacts have been recorded by Google, which experienced transient faults causing incorrect outputs in its production environment [hochschild2021cores]. In 2018, faced with the ongoing risk of transient faults on its large-scale infrastructure, Meta launched an internal investigation to seek solutions [dixit2021silent]. Transient faults may cause either fail-stop errors, which lead to system crashes, or fail-continue errors, which produce incorrect outcomes. While fail-stop errors can often be addressed through checkpoint/restart mechanisms [phillips2005scalable, NEURIPS2019_9015, tao2018improving, tensorflow2015-whitepaper] or algorithmic approaches [chen2009optimal, wu2011fault, hakkarinen2012multilevel, hakkarinen2014fail, chen2008scalable, chen2008extending], fail-continue errors are more problematic as they silently corrupt the state of applications and result in incorrect outputs [mitra2014resilience, cher2014understanding, dongarra2011international, calhoun2017towards, snir2014addressing]. These errors are particularly dangerous in environments where safety is critical [li2017understanding]. This paper focuses on fail-continue errors within the computational logic units, assuming that memory errors and those causing system stops are managed using error-correcting codes and checkpoint/restart techniques. We refer to these issues as soft errors.
@@
Existing fault tolerance methods are feasible for K-means clustering. Taamneh [taamneh2020parallel] proposes a checkpointing strategy, which involves periodically saving the centroids to stable storage during normal operation and restarting from checkpointed centroids in the event of a failure. However, this method cannot detect silent errors and requires recomputation after an error. Dual modular redundancy (DMR) verifies computational correctness by replicating instructions and comparing the results [oh2002error, oh2002control, reis2005swift, yu2009esoftcheck, chen2016simd]. Figure 1 left illustrates the K-means workflow. We find that DMR can protect the memory-bound routine efficiently, such as the centroids update phase in Figure 1 left step 3. It is because the memory latency of loading the data points is so high that all arithmetic operations can be duplicated without introducing an overhead over $1\%$, even for synchronous instructions like atomicAdd. However, DMR failed to protect the compute-bound distance calculation in Figure 1 left step 1 due to the redundant computations. In Figure 1 left step 2, nearest cluster matching is fused within the distance computation kernel to eliminate redundant memory access. Algorithm-based fault tolerance (ABFT) reduces this redundancy by using checksums based on equivalence relationships, lowering the redundancy to $O(1/N)$, where $N$ is the problem size [chen2008algorithm]. Efforts have been made to apply ABFT on GPUs, designing kernel fusion schemes to minimize the overhead of ABFT by hiding it within the memory transactions and computing unit gaps [ding2011matrix, zhai2021ft, kosaian2021arithmetic, wu2023anatomy].
@@
However, existing ABFT fusion strategies lack architecture-aware optimization. Kosaian [kosaian2021arithmetic] proposes an error detection scheme for tensor-core but the error correction requires a time-redundant recomputation. Although error correction is discussed in [zhai2021ft, wu2023anatomy], they only utilize a register-reusing which becomes outdated for Amepere architecture. The register reusing stage is illustrated in the SIMT GEMM part of Figure 1 and the asynchronous copy is presented in the Tensor-core part. Before Ampere architecture, the global-to-share memory transfer passes through the register file explicitly, enabling register reusing for checksum computation. However, the latest asynchronous memory copy enables a new data path from global to shared memory bypassing the register. Without register reusing, checksum computation requires additional expensive memory read, destroying the carefully designed pipeline. Additionally, these fault tolerance strategies consider matrix multiplication individually, with no awareness of the underlying K-means routine.
@@
Furthermore, in the state-of-the-art K-means implementation, cuML suffers from fixed kernel parameters, resulting in sub-optimal and low utilization of peak computing performance over various problem sizes. For example, the K-means in cuML obtain a performance of less than 10% of the peak performance for both single and double precisions. Due to the underlying tall-and-skinny matrix multiplication computation inside the K-means, a code generation strategy and a parameter selection scheme are necessary to improve the K-means performance for various of input shapes and data types.
@@
To address these issues, we propose FT K-means, a high-performance K-means implementation equipped with an algorithm-based fault tolerance scheme that detects and corrects silent data corruptions at computing units on the fly. More specifically, our contributions include the following:
@@
  • Experimental results of single precision and double precision on an NVIDIA A100 server GPU and a Tesla Turing T4 GPU show that FT K-means offers a $10\%-300\%$ improvement compared to the state-of-the-art library cuML. The fault tolerance scheme in FT K-means introduces an average overhead of 11%, even under tens of error injections per minute.
@@ II-A Fault Model
FT K-means is designed to identify and correct errors in computing units that could influence the final results. It operates under the assumption that memory errors are managed by ECC [bird2017neutron] and communication reliability issues are addressed by FT-MPI [fagg2000ft]. For compute errors during runtime, a fault-tolerant strategy is implemented based on the single-event upset (SEU) assumption [binder1975satellite, petersen2013single, binder1975satellite], which posits that only one soft error occurs within each detection and correction interval. This assumption is supported by the low frequency of multiple soft errors in [reis2005swift, zhai2021ft, wu2023ft, ding2011matrix, wu2014ft]. The fault tolerance method involves high reliability in detecting faults with minimal false alarms [turmon2000software]. Specifically, each threadblock randomly selects an element to corrupt by flipping a single bit, either in its 32-bit float representation or 64-bit double representation. A checksum test with a defined threshold $\delta$ then attempts to operate the corrupted computations, followed by an application of an error correction scheme.
@@ II-C Algorithm-Based Fault Tolerance
Soft error protection algorithms aim to identify and rectify errors that can arise in iterative or computationally demanding applications, such as scientific computing [liu2023cusz, huang2023exploring, jian2024cliz, liu2024high]. The development of these algorithms dates back to 1984, with Huang [huang1984algorithm] first specifically designed for matrix-matrix multiplication. The fundamental concept of these algorithms involves encoding matrices $X$ and $Y$ into checksums $X^{c}$ and $Y^{r}$, respectively. This encoding is achieved through the following equations.
@@
where, $i$ represents the step number in the outer-product update of the matrix $D$, with $D_{i}$ indicating the result of each outer-product, namely $X^{\prime}(:,i)\cdot Y^{\prime}(i,:)$. The offline version of the double-checksum approach can correct only one error in the entire computation. In contrast, the online version corrects a single error at each step of the update, allowing it to address multiple errors throughout the entire program. Ding et al. [ding2011matrix] first presented an implementation of the outer product ABFT-GEMM for GPUs. To minimize the memory latency from checksum operations, Zhai et al. developed combined compute kernels for GEMM on AVX-512-enabled CPUs [zhai2021ft, zhai2023ft]. Kosaian and Rashmi [kosaian2021arithmetic] present a warp-level ABFT implementation capable of error detection, but not correction. Shixun et al. [wu2023anatomy] proposed a fully-fused ABFT-GEMM that effectively detects and corrects computational errors. However, the kernel fusion strategy relies on reusing registers during transfers between global and shared memory. This approach does not extend well to GPU architectures post-Turing, e.g. Ampere, due to the introduction of asynchronous copy instructions. Besides matrix multiplication, the ABFT scheme is widely adopted in fast Fourier transforms [liang2017correcting, wu2024turbofft], sorting [li2019ft], iterative methods [chen2013online, tao2016new, chen2014extending], and computer vision [zhao2020algorithm].
@@ III-A3 Kernel Fusion in thread and thread block level
After GEMM, we write the result matrix back to GPU memory and load the matrix again in order to do a row-wise reduction. This greatly increases the amount of data movement and increases storage overhead. In this part, we apply kernel fusion [zhai2023bytetransformer, zhai2023architectural] to accomplish part of the reduction within the GEMM kernel. Firstly, as Figure 2 Step 2 indicates, each thread in the GEMM kernel handles a small submatrix of the result matrix. After the computation of GEMM, we can simply find the minimum column in each row within this submatrix, and write it to shared memory. After all threads in the threadblock finish this step, thread 0 reads all results in shared memory, calculates another row-wise minimum for each row as the final partial answer for this threadblock, and writes it to GPU memory. Assuming each threadblock has size $TM\times TN$, the reduction kernel only needs to load and handle $TN/N$ of data compared to the last part, which obtains a speedup of 1.13x compared to the last part.
@@ III-B Automatic Code Generation
Default naming of kernel parameters A group of kernel parameters in cuML and CUTLASS refers to a set of parameters, threadblock level parameters, warp level parameters, and thread level parameters. Each level is composed of three parameters from each dimension. For example, in warp level, the three parameters are labeled Warp.M, Warp.N, and Warp.K, and they refer to the three dimensions of M, N, and K in GEMM.
@@
Code Generation Strategy Figure 3 demonstrates the method we used in our code generation strategy. The code structure of the CUTLASS GEMM kernel in cuML is on the left side of the figure. Firstly, the epilogue is integrated into the FusedDistanceNN kernel. The FusedDistanceNN kernel handles both GEMM and reduction. And then the kernel is wrapped into GemmKernel. With threadblock, warp, and thread level parameters set, it is then transferred to the next template as FusedDistanceNNGemm. Finally, with all K-means routine sets, the whole cutlassFusedDistanceNN template function works as an interface of the cluster assignment stage. In order to generate a set of feasible kernels with customized parameters, we write a code to test all possible parameters in the search space defined by ourselves, as shown in the code generation part. And for every group of parameters, try it in a demo code. If it can compile and run, which means it is functionally correct. And then we put it in the parameters queue. With all parameters, we then run a code generator to modify the three parts of the original cuML source code, injecting corresponding functions for each parameter in each stage.
@@ III-B1 Kernel Parameters
The kernel parameters used in code generation is not chosen by brute forcing every possible integer in a range for every parameter in the parameter group (a parameter group means: threadblock level, warp level, and thread level parameters). We follow some rules. 1) all parameters must be power of 2 2). Warp.K = Threadblock.K. 3). warp size/thread size is 8 or 16. 4). thread size is fixed for FP32 (16, 8, 4) and FP64 (8, 8, 4) owing to the size of the tensor core.
@@ III-B2 Code Generation Template
The code generation strategy defines the K-means kernel using different tiling parameters. For single precision, 157 kernels with different parameter sets are defined while 145 kernels are defined for double precision. The test workflow illustrated in Figure 3 checks the feasibility of those kernels and performs the benchmark over 64 problem sizes. The benchmark result of different kernels will be employed as the kernel selection criterion. Below, we give a brief explanation of the K-means kernel so that we can switch to the discussion of the fault tolerance part.
@@ V Performance Evaluation
We evaluate our K-means on two NVIDIA GPUs, a Tesla Turing T4 and a 40GB A100-PCIE GPU. The Tesla T4 GPU is connected to a node with two 16-core Intel Xeon Silver 4216 CPUs, whose boost frequency is up to 3.2 GHz. The associated CPU main memory system has a capacity of 512 GB at 2400 MHz. The A100 GPU is connected to a node with one 64-core AMD EPYC 7763 CPU with a boost frequency of 3.5 GHz. We compile programs using CUDA $\mathtt{11.6}$ with the $\mathtt{-O3}$ optimization flag on the Tesla T4 machine, and using CUDA $\mathtt{12.0}$ on the A100 machine. A100 has a peak computational performance of $19.5$ TFLOPS for single precision and $9.7$ TFLOPS for double precision. The memory bandwidth is $1.55$ TB/s. T4 has a peak performance of $8.1$ TFLOPS for single precision and a peak performance of $0.253$ TFLOPS for double precision. The bandwidth of T4 is $320$ GB/s. We first demonstrate the benchmark result between FT K-means without fault tolerance and cuML for FP32 and FP64 on A100. Next, we evaluate the FT K-means under fault tolerance. Then, we benchmark FT K-means under error injections with cuML and Wu’s ABFT. Finally, we present the same performance evaluation on the T4 GPU. All experimental results are averaged over ten trials.